On July 1, 2016, Robert paid $10,000 for a ten-year U.S. Treasury bond with a stated interest rate of 4%, payable annually on July 1. On February 9, 2017, 223 days after purchasing the bond, Robert sold the bond to Alex for $10,050. Which of the following should be reported on Robert's return?

a) $0 of interest income and $50 of short-term capital gain.

b) $156 of interest income and $50 of short-term capital gain.

c) $244 of interest income and $50 of short-term capital gain.

d) $244 of interest income and $194 of short-term capital gain.
Answer #1

Calculate interest income and short term capital gain

Interest income = 10000*4%*223/365 = $244

Short term capital gain = 10050-10000 = $50

So answer is c) $244 of interest income and $50 of short-term capital gain.